OpenStudy (disco619):

Can someone answer this with Explanation? Q] (a) Three numbers are given in the answer space. Write 'L' against the Largest, 'M' against the next largest and 'S' against the Smallest. i) 0.7 million ii) 687,000 iii) eight hundred and four thousand (b) An amount of Money is divided into two parts in the ratio 1:4. Find the smaller part as a percentage of the Whole amount.

OpenStudy (cwrw238):

convert i) and iii) to same form as ii) then the answer will be obvious

OpenStudy (disco619):

[NEWS FLASH] That's what i want to ask. How? xD

OpenStudy (bibby):

Just read iii out loud. 804 thousands = 804,000

OpenStudy (disco619):

What's up with that million? 1 million is 100,000 right? So 0.7 would be?

OpenStudy (cwrw238):

first one = 0.7 * 1,000,000

OpenStudy (cwrw238):

no i million = 1,000,000

OpenStudy (disco619):

Oh...

OpenStudy (cwrw238):

0.1 of 1,000,000 = 100,000 so 0.7 of 1,000,000 = 7 * 100,000

OpenStudy (cwrw238):

0.1 = 1/10

OpenStudy (cwrw238):

1/10 of 1,000,000 = 1,000,000 / 10 = 100,000

OpenStudy (disco619):

Which means 0.7 millions = 700,000???

OpenStudy (disco619):

0.7 million = M 687,000 = S eight hundred and four thousand = L

OpenStudy (disco619):

What about the (b)????

OpenStudy (anonymous):

Two parts in the ratio of 1:4?

OpenStudy (disco619):

(b) An amount of Money is divided into two parts in the ratio 1:4. Find the smaller part as a percentage of the Whole amount.

OpenStudy (disco619):

\[\frac{ 1 }{ 4 } \times 100\] ???

OpenStudy (anonymous):

Sorry, it's just worded confusingly. Here, set it up more like this\[\frac{ 1 }{ 4 } \times \frac{ 100 }{ 1 } = \frac{ 100 \times 1 }{ 4 \times 1 }\]

OpenStudy (disco619):

Yes, it's a Question from an Exam Paper... They have a Habit of Making stuff Confusing >:(

OpenStudy (anonymous):

so, 100 / 4 = 25. Now two parts of that?

OpenStudy (disco619):

I think the Question means that the Money is divided into two parts. One part is 1 and the other 4, making a ratio 1:4. We have to find the percentage of 1

OpenStudy (anonymous):

So, wouldn't we get 1/5 parts then?

OpenStudy (disco619):

Aw yessssss

OpenStudy (anonymous):

Alright, so 100 / 5 Would be our fraction then.

OpenStudy (disco619):

Ok so the Answer is Correct. But can you explain a *little* how would we get 100/5 so there is no confusion. Answer is 20 (correct)

OpenStudy (anonymous):

Alright, \[\frac{ 100 }{ 1 } \times \frac{ 1 }{ 5 } = \frac{ 100 \times 1 }{ 1 \times 5 } = \frac{ 100 }{ 5 }\]

OpenStudy (disco619):

THANKS AGAIN! lol